Use the information in the table to answer the question.
Year 2015 2016 2017 2018
Nominal Salary $53,000 $53,800 $55,000 $57,000
Price Level 120 123 128 131
Real Salary

What was the Average per year % increase in REAL salary between 2015 and 2018 ?

-0.48%

-0.51%

-0.52%

-0.50%
Answer #1

Nominal Growth in Salary = (57,000/53,000)1/3 - 1

Nominal Growth in Salary = 2.46%

Inflation Growth Rate = (131/120)1/3 - 1

Inflation Growth Rate = 2.97%

Real Growth in Salary = 0.0246 - 0.0297

Real Growth in Salary = -0.51%